### v1

# How to verify that a given process is a Lévy-Driven Ornstein-Uhlenbeck Process

Abstract
[…]
Ornstein-Uhlenbeck process,
Lévy process,
Sampled process,
Model verification,
Test statistics,
Sample correlation ,
keywords:
[…]
11footnotetext: Department of Mathematics and Statistics, Rhodes College, Memphis, TN 38112, USA. Emails: abdelrazeqi@rhodes.edu, smirh-25@rhodes.edu, zhadi-25@rhodes.edu
1 Introduction
The continuous time analogue of the AR(1) time series is the Ornstein-Uhlenbeck process (equivalently, CAR(1) or CARMA(1,0) process) $Y$, which is the unique stationary solution of the stochastic differential equation
$dY(t)=-aY(t)dt+\sigma dL(t),~{}a,\sigma>0.$ (1.1)
The driving process $L$ is generally assumed to be Lévy - i.e. $L$ has independent increments. In [4], Barndorff-Nielsen and Shephard used this process in a stochastic volatility model, and it has subsequently received considerable attention in the literature. As Estate Khmaladze points out in [8], scanning through the database MathSciNet with key-word “Ornstein-Uhlenbeck process” returned 18 pages of results, about 360 titles, all dated after year 2000. When the same is searched today, the database returned 68 pages of results, about 1360 titles, all dated after year 2000.
Between 2014 and 2018, Abdelrazeq et al. developed theories and results indicating how to test if a process is a Lévy-driven CAR(1) process [[1], [2], [3]]. Using these results, we develop a step-by-step methodological approach for how these findings can be used to verify the process and test its driving process distribution. The performance of these test statistics under the null hypothesis is illustrated via simulation studies. To demonstrate the power of the test, we computed the power under one alternative case. We then illustrate how to apply the methodology step-by-step to many economic and financial data examples.
We hope that this will encourage many authors to adopt this methodology to verify the model before applying it. The objective of this paper is to make these results, along with the accompanying R code, accessible to users from all fields who may want to fit and apply this model to their data.
[…]
We proceed as follows. In Section 2, we formally introduce the model and the test statistic. In Section 3, we detail our methodology for testing various background processes (Brownian motion, Gamma, Inverse Gaussian, etc.). We refer to Sections 1 and 2 of our supplemental material for a detailed explanation of our methodology, as well as corresponding R code. In Sections 4 and 5, we produce examples that apply our test to high frequency financial data: S&P 500 trading data and Euro/USD currency exchange data, respectively. We refer to Section 3 of our supplemental material for the R code corresponding to the S&P 500 example in Section 4.
2 Preliminaries
2.1 Lévy processes
[…]
Definition 2.1 (Lévy process).
[…]
  • $L(0)=0$ a.s.
[…]
  • $L$ has stationary increments, i.e., $L(t+s)-L(s)$ has the same distribution as $L(t)$, for any $s,t>0$.
  • $L$ is stochastically continuous, i.e. $\forall~{}\epsilon>0$ and $\forall~{}0\leq s<t<\infty$,
    $\lim_{s\to t}P(|L(t)-L(s)|>\epsilon)=0.$
  • $L$ has càdlàg (right continuous with left limits) sample paths.
Definition 2.2.
[Second-Order Lévy Process] We define $L$ to be a second-order Lévy process if $L$ is a Lévy process and $\mathbb{E}\left[L^{2}(1)\right]<\infty$. If $\mu=\mathbb{E}\left[L(1)\right]$ and $\eta^{2}=\mathrm{Var}\left[L(1)\right]$, then by the independence and stationarity of the increments of $L(t)$ we have
$\mathbb{E}\left[L(t)\right]$ $=$ $\mu t,\textbf{{ } { } }t\geq 0,$
$\mathrm{Var}\left(L(t)\right)$ $=$ $\eta^{2}t,\textbf{{ } { } }t\geq 0.$ (2.1)
Examples of second-order Lévy processes include Brownian motion with drift, the Poisson process, Inverse Gaussian process, Beta process, and the Gamma process, which is characterized by
$L(1)\sim\Gamma(\alpha,\beta)=\Gamma\left(\frac{\mu^{2}}{\eta^{2}},\frac{\eta^{2}}{\mu}\right).$
2.2 Lévy-driven CAR(1) models
In what follows, we assume that the process $L$ is càdlàg with stationary increments.
Definition 2.3 (CAR(1) process).
A CAR(1) process $Y=\{Y(t),t\geq 0\}$ driven by the process $L=\{L(t),t\geq 0\}$ is defined to be the solution of the stochastic differential equation
$dY(t)=-aY(t)dt+\sigma dL(t),$ (2.2)
where $a,\sigma\in\mathbb{R}_{+}$ and $Y(0)$ is independent of $\{L(t),t\geq 0\}$. We call the process $L$ the driving process, and if $L$ is a Lévy process then $Y$ is called a Lévy-driven CAR(1) process.
Lemma 2.4 ([4]).
Let $Y$ be a strictly stationary CAR(1) process driven by a second-order Lévy process $L$ such that (2.2) holds. Then for $s\geq 0$,
$\mathbb{E}\left[Y(0)\right]$ $=$ $\frac{\mu\sigma}{a},~{}~{}~{}\gamma_{Y}(s)\equiv\mathrm{Cov}(Y(0),Y(s))=\frac{\sigma^{2}\eta^{2}}{2a}e^{-as}.$ (2.3)
In this study, we consider the Lévy-driven CAR(1) process sampled discretely at intervals of length $1/M$ over a time interval $[0,N]$, hence we have $N\times M$ observations sampled at $1/M,2/M,\ldots,N$. The following equations and Lemma 2.5 have been introduced and proved in [2], and we are including them here to introduce the concepts and provide a smooth thinking process for users.
Let $Y$ be a Lévy-driven CAR(1) process, denote
[…]
to be the increment of the (unobserved) Lévy process $L$ over the unit interval $[n-1,n]$. This unobserved increment, using the trapezoidal approximation as in [5], can be approximated by
$\Delta_{1}\widehat{L}^{(M)}_{n}$$\equiv$ $\frac{a}{M\sigma}\sum_{i=(n-1)M+1}^{nM}Y_{\frac{i}{M}}+\left(\frac{1}{\sigma}-\frac{a}{2M\sigma}\right)\left(Y_{n}-Y_{n-1}\right).$ (2.5)
Let $\overline{\Delta_{1}\widehat{L}^{(M)}}$, $\widehat{\eta}^{2}$ and $\widehat{\gamma_{\Delta_{1}\widehat{L}^{(M)}}}(k)$ be the sample mean, sample variance and sample covariances at lag $k\geq 1$, respectively, of the estimated increments $\Delta_{1}\widehat{L}^{(M)}_{n}$, i.e.
$\overline{\Delta_{1}\widehat{L}^{(M)}}\equiv\frac{1}{N}\sum_{n=1}^{N}\Delta_{1}\widehat{L}^{(M)}_{n},~{}~{}\widehat{\eta}^{2}\equiv\frac{1}{N}\sum_{n=1}^{N}\left(\Delta_{1}\widehat{L}^{(M)}_{n}-\overline{\Delta_{1}\widehat{L}^{(M)}}\right)^{2}$ (2.6)
and
$\widehat{\gamma_{\Delta_{1}\widehat{L}^{(M)}}}(k)\equiv\frac{1}{N-k}\sum_{n=1}^{N-k}\left(\Delta_{1}\widehat{L}^{(M)}_{n+k}-\overline{\Delta_{1}\widehat{L}^{(M)}}\right)\left(\Delta_{1}\widehat{L}^{(M)}_{n}-\overline{\Delta_{1}\widehat{L}^{(M)}}\right).$ (2.7)
Lemma 2.5.
Under the assumptions of Lemma 2.4, we have:
  • (i)
    $\widehat{\gamma_{\Delta_{1}\widehat{L}^{(M)}}}(k)\stackrel{{\scriptstyle p}}{{\longrightarrow}}0~{}~{}\text{as}~{}~{}N\wedge M\to\infty~{}~{}\forall~{}k\geq 1$;
  • (ii)
    $\sqrt{N}~{}\widehat{\gamma_{\Delta_{1}\widehat{L}^{(M)}}}(k)\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,\eta^{4})~{}~{}\text{as}~{}~{}N\to\infty~{}\text{and}~{}N/M\to 0~{}~{}\forall~{}k\geq 1.$
  • (iii)
    $W_{\Delta_{1}\widehat{L}^{(M)}}(k)\equiv\sqrt{N}~{}\frac{\widehat{\gamma_{\Delta_{1}\widehat{L}^{(M)}}}(k)}{\widehat{\eta}^{2}}\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,1)~{}~{}\text{as}~{}~{}N\to\infty~{}\text{and}~{}N/M\to 0.$
2.3 Test Statistic
We define $W_{\widehat{\Delta_{1}\widehat{L}^{(M)}}}(1)$ to be the test statistic $W_{\Delta_{1}\widehat{L}^{(M)}}(k)$ defined in [1] if we replace the parameter $a$ by an estimator $\widehat{a}^{(M)}_{N}$ to be specified in the next section and we define $\widehat{\widehat{\eta}^{2}}$ to be ${\widehat{\eta}^{2}}$ if we replace $a$ by $\widehat{a}^{(M)}_{N}$.
Let $Y_{t}$ be a discretely sampled stochastic process. If $Y$ is a CAR(1) model driven by a process $L$, we can use the estimated increments to test $H_{0}$ that $L$ has uncorrelated increments, which will be true if $L$ is a Lévy process. We reject $H_{0}$ for a large absolute value of the statistic $W_{\widehat{\Delta_{1}\widehat{L}^{(M)}}}(1)$,
where
$W_{\widehat{\Delta_{1}\widehat{L}^{(M)}}}(1)\equiv\sqrt{N}~{}\frac{\widehat{\gamma_{\Delta_{1}\widehat{L}^{(M)}}}(1)}{\widehat{\eta}^{2}}.$ (2.8)
Under $H_{0}$, for large $N,M$ and $\frac{N}{M}$ small we have
$\alpha\approx P\left(|W_{\widehat{\Delta_{1}\widehat{L}^{(M)}}}(1)|>z_{\alpha/2}\right).$ (2.9)
3 Methodology
In this section, we outline the step-by-step methodology for testing different background processes. The approach and results are clearly presented to allow users to apply these techniques to their own models and data. See Sections 1 and 2 of the supplemental material for further explanation the methodology along with corresponding R code.
3.1 Step 1: Choosing the right estimator for $a$
We are concerned with the consistency of the test statistic $W_{\Delta_{1}\widehat{L}^{(M)}}(1)$ defined if we replace the parameter $a$ by an estimator. Thus, we must first choose the most accurate estimator for $a$.
Least squares based estimator
We need to consider an estimator for $a$ for the general second order driving process $L$ (i.e. $\mu\neq 0$). Consider the following estimator of $a$:
$\widehat{{a}}_{N}^{(M)}=\frac{\sum_{n=1}^{NM}\left(Y_{\frac{n-1}{M}}-Y_{\frac{n}{M}}\right)\left(Y_{\frac{n-1}{M}}-\overline{Y}\right)}{\frac{1}{M}\sum_{n=1}^{NM}\left(Y_{\frac{n-1}{M}}-\overline{Y}\right)^{2}}~{}~{}\text{where}~{}~{}\overline{Y}=\frac{1}{NM}\sum_{n=1}^{NM}Y_{\frac{n}{M}}.$ (3.1)
This estimator is based on the least squares approach as in [1]; hence we will refer to it as the least squares based estimator or LSB. This estimator is recommended for use when the driving process is an unspecified Lévy process, or when Brownian motion is expected to be involved in the driving process. Moreover, one should use this estimator in order to test if the driving process is Brownian motion or not. The explanation and R code for this estimator can be found in Section 1.4.1 of the supplemental material.
Davis-McCormick based estimator
Brockwell et al. 2007 [5] introduced an alternative estimator for $a$:
$\widehat{a}_{N}^{(M)}=\sup_{0\leq n<[NM]}\frac{\log(Y_{\frac{n}{M}})-\log(Y_{\frac{n+1}{M}})}{\frac{1}{M}}$ (3.2)
This estimator is based on the highly efficient Davis-McCormick estimator as in [7]; hence we refer to it as Davis-McCormick based estimator or DMB estimator. This estimator is incredibly accurate as compared to the LSB estimator; however, it should not be used if Brownian motion is expected to be involved in the driving process. We have tested and recommend this estimator under Gamma, Inverse Gaussian, Beta, and any mixed combinations of such Lévy processes.
In general, the DMB estimator does not take negative values for $Y$. Hence, if $Y$ is negative at any point, then use the LSB estimator. The explanation and R code for this estimator can be found in Section 1.4.2 of the supplemental material.
Below in Tables 1-3, we observe the comparison of the accuracy of our two estimators in cases when the driving process is not Brownian motion, and we fix $\sigma=1$, $\mu=1$, $\eta=1$, $N=100$, $M=100$, and $\lambda=1$ in relevant cases.
Gamma driven CAR(1) process
Table 1: We fix $\{\sigma=1,\mu=1,\eta=1,N=100,M=100\}$
$a$ LSB DMB
0.3 0.2905178 0.3000090
0.9 1.0248610 0.9000810
5 5.3899950 5.0025020
10 9.6480590 10.010010
For the Gamma driven process, we observe that the DMB estimator is more accurate in estimation for all values of $a$.
Inverse Gaussian driven CAR(1) process
Table 2: We fix $\{\sigma=1,\mu=1,\eta=1,N=100,M=100\}$
$a$ LSB DMB
0.3 0.2882148 0.2998861
0.9 0.7399785 0.8998923
5 5.2126420 5.0017530
10 9.3897690 10.009330
For the Inverse Gaussian driven process, we observe that the DMB estimator is more accurate in estimation for all values of $a$.
Mixed Inverse Gaussian and Gamma driven CAR(1) process
Table 3: We fix $\{\sigma=1,\mu=1,\eta=1,N=100,M=100\}$
$a$ LSB DMB
0.3 0.2879226 0.2999216
0.9 0.8566651 0.8998544
5 4.7511070 5.0017370
10 9.4394340 10.009760
For the mixed Inverse Gaussian and Gamma driven process, we observe that the DMB estimator is more accurate in estimation for all values of $a$.
We observe that the DMB estimator is more precise than the LSB estimator in most cases, yet both are relatively accurate in estimation for each background process. Moreover, the LSB estimator should be used when the driving process is Brownian motion or any unspecified driving process.
3.2 Step 2: Choosing N and M
Before calculating the recovered increments, one should choose $N$ and $M$ such that $N/M$ is small, c.f. [2]. $N$ can represent a day, a month, or a year where $M$ is the number of observations in each period accordingly. Also, we have observed that for small $a$ and $N$, the test has had some issues with the p-value being close to the nominal level. Hence, we recommend that N be more than 50.
3.3 Step 3: Recovering increments with estimated $a$
In order to proceed, we define the recovered increments using an estimator of $a$:
$\widehat{\Delta_{1}\widehat{L}^{(M)}_{n}}$$\equiv$ $\frac{\widehat{a}^{(M)}_{N}}{M}\sum_{i=(n-1)M+1}^{nM}Y_{\frac{i}{M}}+\left(1-\frac{\widehat{a}^{(M)}_{N}}{2M}\right)\left(Y_{n}-Y_{n-1}\right)$
See Section 1.7 of our supplemental material for R code alongside further explanation of the recovered increments with estimated $a$.
3.4 Step 4: Calculating the test statistic
We calculate $W_{\widehat{{\Delta_{1}\widehat{L}^{(M)}}}}(1)$ as defined in Equation (2.8), then for the significance level $\alpha=0.05$, we reject the null hypothesis of uncorrelated increments when $|W_{\widehat{{\Delta_{1}\widehat{L}^{(M)}}}}(1)|>1.96$. In this case, we reject the null hypothesis that the data is Lévy-driven for every background process. If $|W_{\widehat{{\Delta_{1}\widehat{L}^{(M)}}}}(1)|<1.96$, then we fail to reject the null, and it is possible that Lévy-driven CAR(1) is a good model, so afterwards, one may want to check the background process. In this case, we recommend that you continue to Step 5 (see subsection 3.5). Refer to Section 1.8 of the supplemental material for how to calculate the test statistic in R.
Note that a good indication that a test is performing well under the null is that for a significance level of 0.05, and a number of simulations, 400 ($R=400$ in the following Tables), the p-value should be around $0.05\pm 0.02$, which is true in our most cases. It is acceptable to be off one out of twenty, by chance, by the binomial distribution.
Performance of the test for various background processes:
We have defined a Lévy-driven CAR(1) process, a method for recovering the observed increments with an estimated $a$, and the test statistic $W_{\widehat{{\Delta_{1}\widehat{L}^{(M)}}}}(1)$. We now observe the performance of our verification process for various background processes (i.e. Brownian Motion, Gamma, Inverse Gaussian, Beta, and mixed Lévy processes). The tables for each background process display the p-value, see Equation (2.9), for different $N$, $M$, and estimated $a$ values. See Section 2.1 of the supplemental material for the performance of the various background processes along with corresponding R code.
Brownian motion driven CAR(1) process
We use the LSB estimator to estimate $a$ for this process.
Table 4: We fix $\{\sigma=1,\mu=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\alpha}_{\triangle_{1}\hat{BM}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{BM}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{BM}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{BM}^{(M)}}$
$a=0.1$ 0.0225 0.0350 0.0325 0.0475
$a=0.3$ 0.0100 0.0300 0.0125 0.0225
$a=0.5$ 0.0075 0.0050 0.0050 0.0125
$a=0.9$ 0.0200 0.0225 0.0175 0.0175
$a=3$ 0.0475 0.0600 0.0550 0.0275
$a=5$ 0.0375 0.0525 0.0600 0.0450
$a=7$ 0.0400 0.0375 0.0375 0.0550
$a=10$ 0.0375 0.0575 0.0475 0.0425
For the Brownian motion driven process, we observe p-values very near $\alpha=0.05$ for most combination of $N$, $M$, and $a$ values, and it seems it seems that the ratio $N/M$ has no effect. For $a=0.3$, and $a=0.5$, we observe the p-values are close to the nominal level, and we refer to [2] for further information on the effect of the LSB estimator on the result.
Gamma driven CAR(1) process
We use the DMB estimator to estimate $a$ for this process.
Table 5: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\alpha}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{G}^{(M)}}$
$a=0.1$ 0.0375 0.0400 0.0475 0.0275
$a=0.3$ 0.0225 0.0325 0.0350 0.0425
$a=0.5$ 0.0325 0.0525 0.0375 0.0450
$a=0.9$ 0.0325 0.0375 0.0475 0.0350
$a=3$ 0.0525 0.0350 0.0575 0.0475
$a=5$ 0.0325 0.0300 0.0475 0.0475
$a=7$ 0.0350 0.0400 0.0475 0.0425
$a=10$ 0.0300 0.0450 0.0375 0.0425
For the Gamma driven process, we observe p-values very near $\alpha=0.05$ for every combination of $N$, $M$, and $a$ values, and it seems that the ratio $N/M$ has no effect.
Inverse Gaussian driven CAR(1) process
We use the DMB estimator to estimate $a$ for this process.
Table 6: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\alpha}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{IG}^{(M)}}$
$a=0.1$ 0.0375 0.0275 0.0350 0.0525
$a=0.3$ 0.0250 0.0500 0.0575 0.0350
$a=0.5$ 0.0350 0.0400 0.0425 0.0275
$a=0.9$ 0.0500 0.0300 0.0375 0.0300
$a=3$ 0.0325 0.0325 0.0225 0.0275
$a=5$ 0.0400 0.0400 0.0200 0.0200
$a=7$ 0.0275 0.0275 0.0425 0.0150
$a=10$ 0.0400 0.0500 0.0250 0.0350
For the Inverse Gaussian driven process, we observe p-values very near $\alpha=0.05$ for every combination of $N$, $M$, and $a$ values, and it seems that the ratio $N/M$ has no effect.
Mixed Inverse Gaussian and Gamma driven CAR(1) process
Table 7: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\alpha}_{\triangle_{1}\hat{IGG}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{IGG}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{IGG}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{IGG}^{(M)}}$
$a=0.3$ 0.0350 0.0300 0.0225 0.0325
$a=0.9$ 0.0375 0.0375 0.0300 0.0525
$a=5$ 0.0300 0.0350 0.0275 0.0275
$a=10$ 0.0200 0.0425 0.0350 0.0300
For the mixed Inverse Gaussian and Gamma driven process, we observe p-values very near $\alpha=0.05$ for every combination of $N$, $M$, and $a$ values, and it seems that the ratio $N/M$ has no effect.
Our results emphasizes the strength of the verification process, as we fail to reject the null hypothesis of uncorrelated increments for many Lévy-driven background processes and with varying combinations of $N$, $M$, and $a$ values. See Section 2 of the supplemental material for testing the performance of various background processes.
3.5 Step 5: Testing for the driving process
If the null hypothesis that the driving process is Lévy is not rejected, then we want to test whether the driving process is Brownian motion or not. Another interesting test is whether the driving process is some other specific process, say Gamma, Inverse Gaussian, or any other unspecified process. Here we recommend the following two procedures, generally taken from [3] and [10], and modified to our verification process, in order to accomplish these objectives.
Procedure 1:
If we want to test if the driving process is BM or not, then we follow these steps.
  1. 1.
    Recover the estimated increments, as in Equation (3.3), using the LSB estimator.
  2. 2.
    Find a one bootstrap sample from the calculated estimated increments
  3. 3.
    Find the sample mean and standard deviation of the bootstrapped sample found in step 2
  4. 4.
    Conduct KS test with the sample mean and standard deviation calculated in step 3.
This procedure should only be used if the process that you are testing for is Brownian motion, see [6].
Procedure 2:
Taken from [10], if we want to test whether the driving process is some process other than Brownian motion, i.e., Gamma, Inverse Gaussian, or any other unspecified process, we follow these steps:
  1. 1.
    Recover the estimated increments, as in Equation 3.3, using the appropriate estimator (LSB or DMB)
  2. 2.
    Assuming the recovered increments follow a particular distribution $F(.;\theta)$, estimate the parameters for this distribution from the recovered increments
  3. 3.
    Evaluate $Z_{i}$ the CDF values of all recovered increments - using the estimated parameters from Step 2
  4. 4.
    Rearrange $Z_{i}$ in ascending order
  5. 5.
    Calculate the test statistic $D_{N}$ for the $N$ recovered increments $Z_{i}$, where $i$ goes from $1,...,N$, and $D_{N}$ is defined as
    $D_{N}=N^{1/2}\underset{1\leq i\leq N}{\mathrm{max}}\left\{\frac{i}{N}-Z_{i:N},Z_{i:N}-\frac{(i-1)}{N}\right\}$
  6. 6.
    Generate 1000 random bootstrap samples from the assumed $F(.;\theta)$ , using the estimated parameters from Step 2. For each bootstrap sample:
    1. a)
      Estimate the bootstrap parameters as in Step 2
    2. b)
      Evaluate $Z_{i}$ using the estimated bootstrap parameters
    3. c)
      Rearrange $Z_{i}$ in ascending order
    4. d)
      Calculate the test statistic $D_{N}$ using the formula from Step 5
  7. 7.
    Compare the original test statistic from Step 5 to the 95th percentile critical value of the bootstrap distribution. If it exceeds, we reject the null hypothesis
Theoretically, this procedure can be used when you are testing for any specified process completely determined by mean $\mu$ and variance $\eta^{2}$, see [3] for more details. We have only successfully tested it when the driving processes are Gamma, Inverse Gaussian, and Brownian motion, but as said above, we recommend that you use Procedure 1 when testing for Brownian motion, as it is easier computationally.
Under the $H_{0}$, we expect the p-value for test to be around 0.05. In table 8 below, we use Procedure 1 to test if the driving process is Brownian motion or not with varying $N$, $M$, and $a$ values. We refer to Section 2.2.1 of the supplemental material for the corresponding R code.
Brownian motion driven CAR(1) process using Procedure 1:
Table 8: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\alpha}_{\triangle_{1}\hat{BM}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{BM}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{BM}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{BM}^{(M)}}$
$a=0.3$ 0.0525 0.0400 0.0225 0.0575
$a=0.9$ 0.0350 0.0325 0.0400 0.0400
$a=5$ 0.0475 0.0400 0.0325 0.0475
$a=10$ 0.0450 0.0400 0.0250 0.0325
When the driving process is Brownian motion, we observe p-values below or very near 0.05 for each combination of $N$, $M$, and $a$ values when using Procedure 1.
In tables 9 and 10 below, we use Procedure 2 and test whether the driving process is some other specific process, i.e., Gamma in table 9 and Inverse Gaussian in table 10, under the $H_{0}$. We refer to Section 2.2.2 of the supplemental material for the corresponding R code.
Gamma driven CAR(1) process:
Table 9: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\alpha}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{G}^{(M)}}$
$a=0.3$ 0.0375 0.0475 0.0300 0.0425
$a=0.9$ 0.0525 0.0525 0.0725 0.0375
$a=5$ 0.0475 0.0425 0.0475 0.0725
$a=10$ 0.0600 0.0575 0.0550 0.0600
When the driving process is Gamma, we observe p-values below or very near 0.05 for each combination of $N$, $M$, and $a$ values.
Inverse Gaussian driven CAR(1) process:
Table 10: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\alpha}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\alpha}_{\triangle_{1}\hat{IG}^{(M)}}$
$a=0.3$ 0.0800 0.0725 0.0625 0.0500
$a=0.9$ 0.0400 0.0500 0.0650 0.0425
$a=5$ 0.0675 0.0700 0.0475 0.0350
$a=10$ 0.0625 0.0500 0.0475 0.0775
When the driving process is Inverse Gaussian, we observe p-values below or very near 0.05 for each combination of $N$, $M$, and $a$ values.
3.5.1 Power of the test
When the process is not driven by Brownian motion, i.e., Gamma, Inverse Gaussian, and mixed combinations of these, we expect high p-values for the test when testing for Brownian motion. In tables 11-13 below, the results for the power can be observed when using Procedure 1.
Simulating Gamma driven CAR(1) process and testing for BM using Procedure 1:
Table 11: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\beta}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{G}^{(M)}}$
$a=0.3$ 0.4550 0.8350 0.8400 0.8550
$a=0.9$ 0.4600 0.8450 0.9050 0.8725
$a=5$ 0.4500 0.8900 0.9075 0.9000
$a=10$ 0.5200 0.9225 0.8950 0.8750
When the driving process is Gamma, we observe high p-values for each combination of $N$, $M$, and $a$ values.
Simulating Inverse Gaussian driven CAR(1) process and testing for BM using Procedure 1:
Table 12: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\beta}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IG}^{(M)}}$
$a=0.3$ 0.6025 0.9425 0.9425 0.9475
$a=0.9$ 0.6125 0.9700 0.9425 0.9575
$a=5$ 0.6975 0.9425 0.9675 0.9650
$a=10$ 0.6700 0.9675 0.9600 0.9750
When the driving process is Inverse Gaussian, we observe high p-values for each combination of $N$, $M$, and $a$ values.
Simulating mixed Inverse Gaussian and Gamma driven CAR(1) process and testing for BM using Procedure 1:
Table 13: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\beta}_{\triangle_{1}\hat{IGG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IGG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IGG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IGG}^{(M)}}$
$a=0.3$ 0.5450 0.9000 0.9150 0.8975
$a=0.9$ 0.5475 0.8875 0.8975 0.9175
$a=5$ 0.6625 0.9450 0.9400 0.9250
$a=10$ 0.5800 0.9525 0.9200 0.9350
When the driving process is mixed Inverse Gaussian and Gamma, we observe high p-values for each combination of $N$, $M$, and $a$ values.
In order to illustrate the test’s effectiveness using Procedure 2, we observe the p-values for the test when testing for Brownian motion in tables 14-16 below, when using Procedure 2.
Simulating Gamma driven CAR(1) process and testing for BM using Procedure 2:
Table 14: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\beta}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{G}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{G}^{(M)}}$
$a=0.3$ 0.9175 1.0000 1.0000 1.0000
$a=0.9$ 0.9175 1.0000 0.9975 1.0000
$a=5$ 0.9525 1.0000 1.0000 1.0000
$a=10$ 0.9525 1.0000 1.0000 1.0000
When the driving process is Gamma, we observe high p-values for each combination of $N$, $M$, and $a$ values.
Simulating Inverse Gaussian driven CAR(1) process and testing for BM using Procedure 2:
Table 15: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\beta}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IG}^{(M)}}$
$a=0.3$ 0.9775 1.0000 1.0000 1.0000
$a=0.9$ 0.9950 1.0000 1.0000 1.0000
$a=5$ 0.9875 1.0000 1.0000 1.0000
$a=10$ 0.9825 1.0000 1.0000 1.0000
When the driving process is Inverse Gaussian, we observe high p-values for each combination of $N$, $M$, and $a$ values.
Simulating mixed Inverse Gaussian and Gamma driven CAR(1) process and testing for BM using Procedure 2:
Table 16: We fix $\{\sigma=1,\mu=1,\eta=1,R=400\}$
$N=50,M=100$ $N=100,M=100$ $N=100,M=300$ $N=100,M=500$
$\hat{\beta}_{\triangle_{1}\hat{IGG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IGG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IGG}^{(M)}}$ $\hat{\beta}_{\triangle_{1}\hat{IGG}^{(M)}}$
$a=0.3$ 0.9650 1.0000 1.0000 1.0000
$a=0.9$ 0.9850 1.0000 1.0000 1.0000
$a=5$ 0.9600 1.0000 1.0000 1.0000
$a=10$ 0.9725 1.0000 1.0000 1.0000
When the driving process is mixed Inverse Gaussian and Gamma, we observe high p-values for each combination of $N$, $M$, and $a$ values.
As a general comment, we observe a very good power of the test for all cases.
Although we include the above tables for only one case of the power, namely when the driving process is Brownian motion using Procedure 1 and Procedure 2, we have tested many combinations of these processes using Procedure 2. For example, when the driving process is Gamma and we test for Inverse Gaussian, and vise versa, the p-values are all similar to or greater than the rejection rates in the above tables.
We observe results that are consistent with our expectations both under the $H_{0}$, and for the power of the test. For different combinations of $N$, $M$ and $a$ values, we observe p-values around 0.05 under the $H_{0}$, and for the power of the test, we observe high p-values when testing for many different processes. The R code for this can be seen in Section 2.2 of the supplemental material.
4 Example with financial data
4.1 Data and model
We apply our defined verification process to minute-by-minute stock prices of the S&P 500 index constituents from December 2014 until December 2015; the R code for this application can be found in Section 3 of the supplemental material. The S&P 500 is a major stock market index tracking the performance of 500 of the largest publicly traded companies in the United States. It is the benchmark index for large U.S. corporations. Our matrix of data for any given S&P 500 constituent on any given day includes the date given as an integer, the time (minute-by-minute) on a 24 hour clock, and the open price, high price, low, price, and close price from that minute [9].
In order to implement our verification process, we define price spread for two stocks A and B with prices $S_{A}(t)$ and $S_{B}(t)$ as
$Y_{t}=\ln\left(\frac{S_{A}(t)}{S_{A}(0)}\right)-\ln\left(\frac{S_{B}(t)}{S_{B}(0)}\right).$ (4.1)
Spread measures the relative change in the prices of two stocks over time. For modeling the spread dynamics, we refer to the OU process driven by Lévy noise in Equation (2.2). In our example, we calculate the spread dynamics for several large-cap S&P 500 stock constituent pairs in corresponding GICS sectors.
The model of the spread dynamic for the pair Abbot Laboratories (ABT) and Danaher Corporation (DHR) is given in Figure 1:
Figure 1: ABT and DHR spread dynamic
The model of the spread dynamic for the pair Amgen and Pfizer is given in Figure 2:
Figure 2: Amgen and Pfizer spread dynamic
4.2 Verification
We follow our methodology detailed in Section 3. Following Step 1 (see Subsection 3.1), we choose our estimator for $a$. We note that price spread takes on negative values, so we use the LSB estimator to estimate $a$. Following Step 2 (see Subsection 3.2), we choose values for $N$ and $M$. The time period is given to us from the data, namely the observed spread dynamic from December 2014 - December 2015. Thus, we set $M=500$ such that the spread dynamic is sampled at $1/M,2/M,...,N$. Following Step 3 (see Subsection 3.3), we recover the spread dynamic with our chosen $M$ and our estimated $a$ values. Following Step 4 (see Subsection 3.4), we calculate the test statistic and test the correlation of our recovered increments. The application of this methodology on our data can be found in Section 3 of the supplemental material.
We model the spread of several combinations of large-cap stocks in corresponding Global Industry Classification Standard sectors to verify that they are Lévy-driven, having test statistics below the critical value $z_{1-\alpha/2}=1.96$.
We observe varying and indefinite results from our test. An example measuring the price spread between Abbot Laboratories (ABT) and Danaher Corporation (DHR) fails to reject the null hypothesis of uncorrelated increments, producing a test statistic of 1.05. In alignment with this, the spread of Apple (AAPL) and Google (GOOG) fails to reject with a test statistic of 1.85, which suggests that a Lévy driven OU model may be a good model for these spreads. In Figures 3 and 4 below, we plot the recovered increments for the spread of ABT and DHR as a time series and as residuals respectively.
Figure 3: Recovered increments of ABT and DHR spread dynamic as a time series
Figure 4: Recovered increments of ABT and DHR spread dynamic as points
We observe that the recovered increments of the ABT and DHR spread dynamic exhibit relatively stable, uncorrelated behavior, suggesting that the Lévy-driven CAR(1) model could be a good fit. And when plotted as points, the residuals of the recovered increments appear to lack significant trends, outliers, or unusual clustering, further supporting the hypothesis of uncorrelated increments for this spread dynamic.
In Figure 5 below, we plot the sample autocorrelation function for the recovered ABT and DHR spread dynamic.
Figure 5: The sample autocorrelation function for the recovered increments of ABT and DHR spread dynamic
With a test statistic of 1.05, which is well below the critical value, $z=1.96$, we fail to reject the null hypothesis, which suggest that a Lévy-driven CAR(1) model could be a good fit for this spread dynamic. This is confirmed by the autocorrelation function being well below the dashed line at lag $=$ 1, so we move to Step 5.
Following Step 5 (see Subsection 3.5), we applied our test to assess whether the driving process of these spreads was Brownian motion or not. Following Procedure 1, we found that the spread dynamic of ABT and DHR failed to reject the normality assumption of the recovered increments, suggesting that the increments are consistent with the Brownian motion driven process. We observe consistent results when we apply Steps 1-5 to the spread of Apple (AAPL) and Google (GOOG). Moreover, we apply Procedure 2 to the spread of ABT and DHR testing for Brownian motion, and observe that we spread dynamic still fails to reject the normality assumption under this procedure. Our applications of Procedure 1 and 2 to our data can be accessed in Sections 3.8.1 and 3.8.2 of the supplemental material.
On the other hand, the spread of Goldman Sachs Group (GS) and Morgan Stanley (MS) reject the null, producing a test statistic of 4.30; moreover, the spread of Amgen Inc (AMGN) and Pfizer (PFE) rejected with a test statistic of 4.78, which suggests that a Lévy driven OU model may not be a good model for these spreads. In Figures 6 and 7 below, we plot the recovered increments for the spread of Amgen and Pfizer as a time series and as residuals, respectively.
Figure 6: Recovered increments of Amgen and Pfizer spread dynamic as a time series
Figure 7: Recovered increments of Amgen and Pfizer spread dynamic as points
We observe that there is a very clear leftover relationship in the recovered increments for the Amgen and Pfizer spread dynamic, which suggesting that they are correlated; that suggestion is also supported by the autocorrelation function in Figure 8. The residuals plot reflects clear trends and clustering, indicating potential correlations between increments that do not align with the uncorrelated assumption of a Lévy-driven CAR(1) process.
In Figure 8 below, we plot the sample autocorrelation function for the recovered Amgen and Pfizer spread dynamic.
Figure 8: The sample autocorrelation function for the recovered increments of Amgen and Pfizer spread dynamic
With a test statistic of 4.78, which is well above the critical value, $z=1.96$, we reject the null hypothesis, which suggest that a Lévy-driven CAR(1) model could be a good fit for this spread dynamic. This is confirmed by the autocorrelation function being well above the dashed line at lag $=$ 1, and we do not need to move on to Step 5.
Results vary largely by GICS sector, but the verification process shows that the Lévy-driven CAR(1) model may not be a good fit for every spread process, including some individual pair combinations as well as several GICS sectors. However, many individual pair combinations as well as several GICS sectors appear to perform well under the verification process, justifying the selection of a Lévy-driven model for their spread processes. Refer to Section 3 of the supplemental material for R code of our spread calculation and use of the verification process and methodology on our data.
5 Example with currency exchange data
5.1 Data and model
We go on to apply our verification process to realized volatility calculated from Euro/USD currency exchange rate data. This data was taken from Olsen Data, ranging 10 years from June 2007 through June 2017, and contains the Euro/USD bid and ask prices recorded at 5-minute intervals. In the Foreign Exchange Market, the bid price is what a currency dealer is willing to pay for a currency, while the ask price is the rate at which a dealer will sell that currency. We clean out missing data, weekends, fixed holidays, and similar calendar effects, and we are left with 2,463 full days of 5-minute interval data–about 709,555 data entries.
With our cleaned data set, we calculate daily returns, find the corresponding realized variance, and then compute the square root to find daily realized stochastic volatility.
Realized volatility is represented by the following equation:
$RV=\sqrt{\sum_{i=1}^{N}r_{t}^{2}},$ (5.1)
where $r_{t}$ is the daily return at time $t$. Here, realized volatility measures the variation in returns for the Euro/USD currency exchange rate.
Below in Figure 9, we model the Euro/USD returns at each given 5-minute interval which we calculated from our data:
Figure 9: Euro/USD 5-minute returns
We then follow Equation 5.1 in order to calculate realized volatility. We model the daily realized volatility below in Figure 10:
Figure 10: Euro/USD daily realized volatility
5.2 Verification
We then follow our methodology detailed in Section 3 to verify whether the realized volatility is a Lévy-driven CAR(1) process. Following Step 1, we choose the DMB estimator to be our estimator for $a$, as realized volatility only takes on positive values and exhibits characteristics consistent with the Gamma driven and Inverse Gaussian driven processes. Following Step 2, we choose the number of large periods to be $N=35$, and the sampling frequency to be $M=70$, where $2463/M=N$; 2463 being the total number of days in our cleaned data set. Following Step 3, we recover the increments for the Euro/USD realized volatility with our chosen values and estimated parameters. Then following Step 4, we calculate the test statistic and test the correlation of the recovered increments.
In Figures 11 and 12 below, we model the recovered increments of the Euro/USD realized volatility both as a time series and as residuals.
Figure 11: Recovered increments of Euro/USD realized volatility as a time series
Figure 12: Recovered increments of Euro/USD realized volatility as points
We observe that the recovered increments of the Euro/USD realized volatility exhibit noticeable correlation over the time period, suggesting that the increments are not independent. When plotted as individual points, the residuals show a noticeable pattern indicating that the increments may exhibit correlation, which suggest that a Lévy-driven CAR(1) model may not be a good fit; that suggestion is supported by the autcorrelation function in Figure 13.
In Figure 13 below, we plot the sample autocorrelation function for the recovered Euro/USD realized volatility.
Figure 13: The sample autocorrelation function for the recovered increments of Euro/USD realized volatility
With a test statistic of 3.76 for the Euro/USD realized volatility, which is well above the critical value, $z=1.96$, we reject the null hypothesis, which suggest that a Lévy-driven CAR(1) model could be a good fit for this spread dynamic. This is confirmed by the autocorrelation function being well above the dashed line at lag $=$ 1, and we do not need to move on to Step 5.
As the test rejects the Lévy-driven CAR(1) model, we do not need to go on to Step 5. However, in alignment with [5], we believe that the Lévy-driven CARMA(2,1) process could be a good fit for realized volatility.
6 Conclusion
In conclusion, our verification process proves very useful in testing the fit of models assumed to be Lévy-driven. In this paper, we test a Lévy-driven model for price spread between S&P 500 stock pairs, as well as Euro/USD exchange data, but the verification process can be applied to other real world data, e.g., interest rates, credit risk, population dynamics, and more. Our test yields strong results, failing to reject the null hypothesis of uncorrelated increments in most cases for many simulated Lévy-driven CAR(1) background processes, i.e., Brownian motion, Gamma, Beta, Inverse Gaussian, and mixed combinations of these processes. Moreover, we observed that when testing for Brownian motion, Procedure 1 and Procedure 2 rejected at high rates when the recovered increments were different from the normal distribution for varying combinations of $N$, $M$, and $a$, and the test rejected at p-values near 0.05 under the $H_{0}$. Our methodology proves a powerful tool in verifying whether a given process is an Ornstein-Uhlenbeck process and has direct application to real world data. This paper details the precise methodology to apply this test, and our supplemental material includes further explanation as well as R code that corresponds with the verification process methodology and our example with financial data. Going forward in our future work, we plan to write a similar, user accessible, methodological paper detailing how to verify that a given process is a Lévy-driven CARMA(2,1) process.